OpenStudy (anonymous):

Copy and complete the table on the previous page into a Word document. Title your journal "Pythagorean Theorem". Discuss any relationships you see in the table? Then, look at the equation below and calculate whether a2 + b2 is >, <, or = to c2 based on the numbers you recorded in your table. a^2+b^2=c^2 a. > b. < c. = Were you able to answer correctly? Why or why not? Please help! Thank you!

OpenStudy (anonymous):

woah ur math question just went from 5th grade math to 7th grade math

OpenStudy (anonymous):

anyways lets answer this :P

OpenStudy (anonymous):

I'm in 8th grade but math is not my strongest point

OpenStudy (anonymous):

can u take a picture of the table plz?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

I took a picture of the whole assignment since it wouldn't let me save the table

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i still need to see a picture of the table

OpenStudy (anonymous):

a^2 + b^2 = c^2 3^2 + 4^2 = c^2 9 + 16 = c^2 √25 = √c^2 c = 5

OpenStudy (anonymous):

ok hold up

OpenStudy (anonymous):

Well first off, this is using the Pythagorean Theorem Method and all the variables are to the power of two. this can be your relationship in the table thing

OpenStudy (anonymous):

Then, look at the equation below and calculate whether a2 + b2 is >, <, or = to c2 based on the numbers you recorded in your table. It basically already tells you that a^2+b^2=c^2

OpenStudy (anonymous):

For the last question, Yes, I was able to answer the question correctly since the table provided me with all the necessary information.

OpenStudy (anonymous):

Thanks again you're a lifesaver!

OpenStudy (anonymous):

No problem :3